OpenStudy (rcp031):

The sum of three numbers is 98. If the ratio of the first to second is 2 :3 and that of the second to the third is 5 : 8, then the second number is

OpenStudy (kainui):

Give the numbers names, like a, b, and c. Then put them in equations as you go.

OpenStudy (mathmate):

or combine the ratios A B C 2:3 *5 5:8 *3 ------ 10:15:24 Once you have a combined ratio, you can figure the individual values by proportion. Example: 4 numbers are in the ratio 2:3:1:4 and the sum of the 4 numbers is 100. Find the four numbers. 1. Calculate the sum 2+3+1+4=10 2. The corresponding numbers are 100*2/10=20 100*3/10=30 100*1/10=10 100*4/10=40 total =100 ok.

OpenStudy (rcp031):

Thnx alot guys you r all amazing mathematicians.